# Manh Khoi Duong 11 $0$-0002-4653-7685 Stefan Conrad

### v2

@@ Abstract
In this paper, we deal with bias mitigation techniques that remove specific data points from the training set to aim for a fair representation of the population in that set. Machine learning models are trained on these pre-processed datasets, and their predictions are expected to be fair. However, such approaches may exclude relevant data, making the attained subsets less trustworthy for further usage. To enhance the trustworthiness of prior methods, we propose additional requirements and objectives that the subsets must fulfill in addition to fairness: (1) group coverage, and (2) minimal data loss. While removing entire groups may improve the measured fairness, this practice is very problematic as failing to represent every group cannot be considered fair. In our second concern, we advocate for the retention of data while minimizing discrimination. By introducing a multi-objective optimization problem that considers fairness and data loss, we propose a methodology to find Pareto-optimal solutions that balance these objectives. By identifying such solutions, users can make informed decisions about the trade-off between fairness and data quality and select the most suitable subset for their application.
@@ 2 Related Work
Drawing inspiration from prior research that has addressed group representativeness [7, 19, 4], we extend the methodology of Duong et al. [9] accordingly. There are many ways to define representativeness. Stoyanovich et al. [19] mentioned two main definitions: proportional representation and coverage. Proportional representation ensures that the dataset contains a representative number of data points from each group. Coverage, on the other hand, ensures that the dataset covers the entire population. We aim for coverage in our work and argue that aiming for proportional representation is not feasible because the group counts are given beforehand and not much can be done about it during the data removal process. The satisfaction of proportional representation also highly depends on user-given constraints for the group sizes. Coverage only requires each group to be represented at least once in the dataset and is hence a more relaxed constraint that does not require any user specification.
@@ 4.2 Data Loss
where $\mathcal{D}_{\text{fair}}\subseteq\mathcal{D}$. A lower value indicates less data was removed, and therefore is better.
@@ 6 Heuristics
To make the framework flexible and agnostic to the fairness metric, we have to use heuristics that only require function evaluations. For this, we use genetic algorithms to solve the optimization problems in Equations (4) and (6). For the multi-objective optimization problem, we use our own modified version of the NSGA-II [6] algorithm. To solve the single-objective problem, we use the genetic algorithm borrowed from Duong et al. [9].
@@ 6.1.1 Random Initializer
as $n$ tends to infinity. This also means that each subset is expected to be half the size of the original dataset.
@@ 7.0.3 Trials
Table 5: Hyperparameter Optimization Results. Best results are in bold.
Dataset Adult Bank Compas
Initializer Selection Crossover Mutation Hypervolume Hypervolume Hypervolume
@@ 7.2.5 Results
Figure 1: Results on the test set using both, single and multi-objective optimization, approaches with varying $\beta$ parameters. The x-axis represents the discrimination value and the y-axis represents the classifiers’ performances. We compare the pre-processed (fair) data with the original data.
@@ 8.1 Data Quality
We extend the work of Duong et al. [9] by introducing new constraints and objectives that the resulting dataset must fulfill. To our knowledge, we are the first to specifically aim for data quality while pre-processing datasets for fairness. Some techniques in the literature [2, 1] violate several data integrity constraints and transform the dataset in a way that makes it unusable. For example, CorrelationRemover in Fairlearn [2] projects discrete features into continuous features. Categorical features such as the label are also affected by it, making it impossible to train classifiers for comparative purposes. The pre-processed datasets from our work are more interpretable and can be used for any further analysis.
@@ 8.2 Trade-off
We proposed a multi-objective optimization problem where the objectives are fairness and data loss. Solving this problem results in a set of Pareto-optimal solutions where each solution is a fair subset. The user can then choose the most suitable subset for their application using the weighted sum of the objectives. Here, we proposed a parameter $\beta$ that weights the fairness objective. The choice of $\beta$ has to be made by the user and depends on the size of the dataset and the importance of fairness in the application. Our suggestions of $\beta$ only serve as a starting point.

### v3

@@ Abstract
In this paper, we deal with bias mitigation techniques that remove specific data points from the training set to aim for a fair representation of the population in that set. Machine learning models are trained on these pre-processed datasets, and their predictions are expected to be fair. However, such approaches may exclude relevant data, making the attained subsets less trustworthy for further usage. To enhance the trustworthiness of prior methods, we propose additional requirements and objectives that the subsets must fulfill in addition to fairness: (1) group coverage, and (2) minimal data loss. While removing entire groups may improve the measured fairness, this practice is very problematic as failing to represent every group cannot be considered fair. In our second concern, we advocate for the retention of data while minimizing discrimination. By introducing a multi-objective optimization problem that considers fairness and data loss, we propose a methodology to find Pareto-optimal solutions that balance these objectives. By identifying such solutions, users can make informed decisions about the trade-off between fairness and data quality and select the most suitable subset for their application. Our method is distributed as a Python package via PyPI under the name FairDo [^1] https://github.com/mkduong-ai/fairdo.
@@ 2 Related Work
Drawing inspiration from prior research that has addressed group representativeness [7, 19, 4], we extend the methodology of Duong et al. [9] accordingly. There are many ways to define representativeness. The work of Stoyanovich et al. [19] explores two main definitions: proportional representation and coverage. Proportional representation ensures that the dataset contains a representative number of data points from each group. Coverage, on the other hand, ensures that the dataset covers the entire population. We aim for coverage in our work and argue that aiming for proportional representation is not feasible because the group counts are given beforehand and not much can be done about it during the data removal process. Coverage only requires each group to be represented at least once in the dataset and is hence a more relaxed constraint.
@@ 4.2 Data Loss
where $\mathcal{D}_{\text{fair}}\subseteq\mathcal{D}$. A lower value indicates less data is removed, and therefore is better.
@@ 6 Heuristics
To make the framework flexible and agnostic to the fairness metric, we have to use heuristics that only require function evaluations. For this, we use genetic algorithms to solve the optimization problems in Equations (4) and (6). To solve the single-objective problem, we use the genetic algorithm borrowed from Duong et al. [9]. For the multi-objective optimization problem, we use our own modified version of the NSGA-II [6] algorithm as described in the following.
@@ 6.1.1 Random Initializer
as $n$ approaches infinity. This implies that each subset is expected to be half the size of the original dataset.
@@ 7.0.3 Trials
Table 5: Hyperparameter optimization results showing hypervolume indicator values for different genetic operators. Best results are highlighted in bold.
Initializer Selection Crossover Mutation Adult Bank Compas
@@ 7.2.5 Results
Figure 1: Results on the test set using two approaches with varying $\beta$ parameters. x-axis and y-axis represent discrimination and performance metrics, respectively.
@@ 8 Discussion
In this section, we reflect on key aspects of our approach, particularly focusing on data quality and the trade-off between fairness and data loss.
@@ 8.1 Data Quality
We extend the work of Duong et al. [9] by introducing new constraints and objectives that the resulting dataset must fulfill. To our knowledge, we are the first to specifically aim for data quality while pre-processing datasets for fairness. Some techniques in the literature [2, 1] violate several data integrity constraints and transform the dataset in a way that makes it unusable. For example, CorrelationRemover in Fairlearn [2] projects discrete features into continuous features. Categorical features such as the label are also affected by it, making it impossible to train classifiers for comparative purposes. The pre-processed datasets from our work do not come with these issues, making them more useful in practice.
@@ 8.2 Trade-off
We proposed a multi-objective optimization problem where the objectives are fairness and data loss. Solving this problem results in a set of Pareto-optimal solutions where each solution is a fair subset. The user can then choose the most suitable subset for their application using the weighted sum of the objectives. Here, we proposed a parameter $\beta$ that weights the fairness objective. The choice of $\beta$ has to be made by the user and depends on the size of the dataset and the importance of fairness in the application. Our suggested values for $\beta$ serve as initial guidance, allowing users to further adjust based on their needs.